# From time crystals to time quasicrystals: Exploring novel phases in transverse field Ising chains

### v1

@@ Abstract
Time crystals (TCs) and time quasicrystals (TQCs) represent novel phases of matter that arise from the breaking of time translational symmetry in periodically and quasiperiodically driven quantum systems. In this study, we explore the formation of a TQC phase within the disordered quantum Ising chain model under a transverse field (ITF). Notably, TQCs demonstrate robust subharmonic responses at multiple incommensurate frequencies, unlike traditional TCs which respond at a single frequency. Our analysis reveals that the TQC phase exhibits stable magnetization responses even in the presence of interaction perturbations and imperfections in the quasiperiodic driving fields. Employing exact diagonalization techniques, we find that increasing the chain length further stabilizes both TC and TQC phases. These results suggest promising pathways for experimental realization of TQCs in cold atomic systems and quantum simulators, opening avenues for deeper investigation into these intriguing dynamical phenomena.
@@ I Introduction
Time crystals (TCs) represent a fascinating phase of matter that emerges from the spontaneous breaking of time translational symmetry (TTS) in periodically driven quantum systemsr1 ; r2 ; r3 ; r4 ; r5 ; r6 . These systems are typically subjected to external periodic perturbations, allowing them to maintain coherence and exhibit long-lived states characterized by discrete time translational symmetry breaking. In this context, preventing Floquet heating is crucial to avoid a transition to infinite temperature, which would lead to the loss of any temporal orderr7 ; r8 .
@@
In disordered closed quantum systems, many-body localization (MBL) acts as a stabilizing mechanism against Floquet heatingr99 ; r9 . MBL enables the system to remain localized in a specific region of the Hilbert space, thus invalidating the eigenstate thermalization hypothesis (ETH). As a result, the quantum order is preserved over extended time scales, as the system retains its initial state within its eigenstates, shielded from thermalization processesr10 ; r11 ; r12 ; Pr12 ; rr12 ; rp12 .
@@
While periodically driven Floquet systems typically exhibit non-conservation of energy due to continuous energy absorption from external fields, the presence of Floquet MBL alters this dynamic. By breaking the ETH, Floquet MBL prevents the system from reaching thermal equilibrium, leading to the emergence of nonequilibrium phases, including time crystalline phasesr13 ; eqr13 ; r14 ; r15 .
@@
In TC phases, the dynamics of observables in the system are characterized by periodic responses that differ from the driving period. Specifically, for a Hamiltonian that oscillates with a period $T$, the response functions of the system will exhibit oscillations with a period $nT$ (where $n\geq 2$)r8 .
@@
Time quasicrystals (TQCs) represent a novel category of dynamical systems distinguished by their capacity to exhibit subharmonic responses to external driving forces, characterized by multiple incommensurate frequencies. Unlike traditional TCs, which are constrained to a singular frequency response, TQCs arise from quasiperiodic driving of a system where the time-dependent Hamiltonian $H(t)$ incorporates irrationally related frequencies. This leads to quasiperiodic dynamics, resulting in a non-repetitive evolution of the system’s state that densely explores a multi-dimensional toroidal phase space, without revisiting previous configurations. Such quasiperiodicity facilitates intricate dynamical behavior, including complex energy level spacings and dense trajectories in phase spacerr17 ; r31 ; r32 .
In contrast to conventional TCs, TQCs exhibit robust subharmonic responses at a minimum of two incommensurate frequenciesSelf-ordered . Specifically, in the TQC phase, some of the order parameters reveals a stable response characterized by multiple frequencies within the quasiperiodically driven system, a phenomenon referred to as the discrete TQC phase. The dynamics of the TQC phase can be interpreted as trajectories traversing a higher-dimensional space with extended periodicityr34 .
In this study, we explore the spontaneous emergence of TQCs within the framework of the quasiperiodically driven quantum Ising chain in a transverse field (ITF) model. In the context of the ITF model, the TQC phase manifests as a robust response in magnetization that reflects multiple frequencies. This phase is characterized by its remarkable stability and coherence in the presence of interaction perturbations and imperfections in quasiperiodic driving fields.
To date, the TQC phases have been observed in various modelsr17 ; FTQ ; FTS ; rr15 ; r18 . In this work, we examine disordered ITF model, initially analyze it under a periodic driving magnetic field, demonstrating the existence of a TC in this system. We further assess the stability of the TC against perturbations from external interactions and imperfections in rotations, finding it to be robust against such disturbances. In the latter part of our study, we investigate the ITF chain under a sinusoidal driving field, establishing that a disordered ITF chain can also exhibit a TQC phase that remains stable against external interactions and rotational imperfections. Our investigation employs exact diagonalization, and our findings indicate that increasing chain length enhances the stability of both TC and TQC phases. Investigating TQCs paves the way for potential experimental realizations in cold atomic systems, quantum simulators, and other platforms capable of simulating quasiperiodic driving.
@@
This paper is structured as follows: In Section II, we introduce the ITF spin chain, which is periodically driven by an external magnetic field, and analyze its magnetization in various directions, demonstrating the existence of a TC phase. We further investigate the robustness of this TC phase against perturbations in Section III. Section IV provides an overview of the definition and general properties of a TQC phase. In Section V, we incorporate a sinusoidal driving field into the disordered ITF chain, revealing the existence of a TQC phase and exploring its stability under different perturbations. Section VI discusses the emergence of a TQC phase within an ITF chain exhibiting a spin glass phase. Finally, we summarize our findings in Section VII.
@@ II TC in disordered ITF model
$H(t)=\begin{cases}\hat{H}_{1}=\sum\limits_{i}\hat{\sigma}_{i}^{x},\quad 0\leq t<t_{1}\\ \hat{H}_{2}=\sum\limits_{i}(-J_{i}\hat{\sigma}_{i}^{z}\hat{\sigma}_{i+1}^{z}+h_{i}\hat{\sigma}_{i}^{x}),\quad t_{1}<t\leq T\end{cases}$ (1)
[…]
@@
The situation is different for finite size chains. The system can tunnel between the two degenerate eigenstates at finite size, and consequently the two lowest lying eigenstates of the system are the following $Z_{2}$ symmetric states: |±$|\pm\rangle=\frac{1}{\sqrt{2}}[|\uparrow\uparrow\dots\uparrow\rangle\pm|\downarrow\downarrow\dots\downarrow\rangle]$. These states are the eigenstates of the Ising parity operator $\hat{P}$ with the eigenvalues ±$\pm 1$, they are long-range correlated with vanishing magnetizationr7 .
At finite temperatures, the Mermin-Wagner theorem asserts that one-dimensional systems with discrete symmetries cannot undergo spontaneous symmetry breaking, as thermal fluctuations in these systems are strong enough to disrupt long-range order, resulting in a finite density of excitationsmermin-wagner ,However, the key point to note is that this theory applies exclusively to equilibrium systems.
@@
In the FM phase, these excitations manifest as domain walls. The energy associated with these domain walls is typically independent of their position along the chain, allowing them to move freely. This mobility of domain walls contributes to the destabilization of any potential long-range order, thus reinforcing the conclusion that spontaneous symmetry breaking cannot be realized in one-dimensional systems at finite temperatures.
@@
The excitation states can be described as superpositions of states featuring finite densities of domain walls distributed across various spatial locations. These delocalized and fluctuating domain walls introduce perturbations that disrupt long-range spin correlations at any finite temperature.
@@
The introduction of disorder, characterized by a non-clean regime with $h_{i}\in[-h,h]$ and $J_{i}\in[J/2,3J/2]$, leads to significant alterations in the system’s physical properties. In the absence of external magnetic fields, the system resides in a FM phase characterized by a non-zero magnetization $m^{z}\neq 0$, as all exchange couplings are positive. Upon the application of a weak magnetic field, specifically when $h\ll J$, domain walls emerge along the chain. For a single domain wall located at position $d$, the excited state of the system is represented as $|d\rangle=|\uparrow\uparrow\dots\uparrow\underline{\downarrow}\uparrow\uparrow\dots\rangle$. In contrast to the clean system, where a thermally activated domain wall can move freely, in the presence of disorder, the domain wall becomes immobilized due to the energy cost associated with its dispersion throughout the chain. As the magnetic field increases, the system can sustain a finite density of domain walls, all of which remain pinned by the disorder. It has been demonstrated that for $h<J$, the system exhibits characteristics of the many-body localized (MBL) phase within the FM phase, which remains stable in the presence of perturbations such as interactionsr23 .
@@
We now examine the potential for spontaneous breaking of time translational symmetry within the phase diagram of the floquet ising Transverse field (FITF) model. The time evolution of the system, as described by equation (1), is governed by the Floquet operator given by:
@@
With $t_{1}=\frac{\pi}{2}$, and $t_{2}=1$, the Floquet operator is simplified to:
@@
Here, the first term captures the time evolution over the interval $\frac{\pi}{2}<t\leq T$ under the Hamiltonian $\hat{H}_{2}$, while the second term represents a spin rotation about the $x$-axis. Through this formalism, we aim to determine the conditions under which time translational symmetry may undergo spontaneous breaking in the context of the FITF model.
Consider the system initially in a fully polarized state along the $z$ axis, represented as $|\psi_{0}\rangle=\otimes_{i=1}^{L}|\uparrow_{i}\rangle$, where $L$ is the total number of spins. By applying the Floquet time evolution operator to this state, we can derive the system’s state at any time $t$, along with the time dependence of various order parameters. We investigate the time dependence of the order parameters, specifically the magnetizations along the $x$, $y$, and $z$ axes, defined as $m^{x,y,z}=\langle\hat{\sigma}^{x,y,z}\rangle$. In the presence of disorder, it is essential to compute ensemble averages of the magnetizations across various random realizations. By averaging over 150 random realizations, we derive the magnetizations $m^{x,y,z}$ for 160 Floquet cycles. Our findings indicate that when $\bar{J}\gg\bar{h}$, the magnetization $m^{z}$ exhibits oscillatory behavior around zero with a period of $2T$ (refer to Fig. 1-(up)). The periodic oscillations of $m^{z}$, with a period that is double that of the Hamiltonian, suggest a breakdown of time translational symmetry. This phenomenon is indicative of time crystal (TC), provided that the observed oscillations remain stable in the thermodynamic limit ($L\rightarrow\infty$) and are resilient against perturbations such as interactions.
Figure 1: Time evolution of the magnetizations along $x,y$, and $z$ directions for a chain of 12 sites. For $J=5.5$ and $h=0.3$ (up), the magnetization along the $z$ axis exhibits periodic oscillations with a period of $2T$, while the $x$ and $y$ components remain zero. The inset plot illustrates the fast Fourier transform (FFT) of $m^{z}(t)$, highlighting a pronounced peak at $\omega=\pi$, which signifies the periodic nature of the magnetization. Conversely, for $J=0.3$ and $h=4.5$ (down), the magnetization oscillations are characterized by a lack of periodicity and stability. The peak value in the FFT is multiplied by 0.01, and this adjustment is applied to all other plots as well.
@@
III Stability of the TC phase
In this section, we investigate the stability of the TC in the presence of various external perturbations. We will demonstrate that the introduction of diverse perturbations, including interactions and imperfect rotations (deviations from perfect rotation), does not compromise the time crystalline order.
III.1 Stability of the TC phase under interactions
In this part, we examine the stability of the TC phase under the influence of exchange interactions along the $x$-axis. In this context, the second term of the Hamiltonian in Eq. (1) is modified as follows:

### v2

# From time crystals to time quasicrystals: Exploring quasiperiodic phases in transverse field Ising chains

@@ Abstract
Time quasicrystals (TQCs) represent a compelling extension of the concept of time crystals (TCs). While TCs break discrete time-translation symmetry by exhibiting a periodic response at a subharmonic of the driving frequency, TQCs display a more complex temporal order. They respond at multiple incommensurate frequencies, values that are not integer multiples of the fundamental driving frequency, resulting in quasiperiodic dynamics. In this work, we investigate the emergence of a TQC in a disordered quantum Ising chain subjected to a quasiperiodic transverse field. Using exact diagonalization, we find that the transverse magnetization exhibits quasiperiodic oscillations which persist over extended prethermal timescales before eventual decay. This indicates that the TQC exists as a long-lived, prethermal dynamical phase rather than a true equilibrium state. We further assess the robustness of this prethermal TQC against interactions and driving imperfections, confirming its stability under realistic experimental conditions. Finite-size analysis reveals that the prethermal TQC lifetime exhibits minimal dependence on system size. Additionally, we explore the emergence of TQCs in the same chain under symmetric sampling of exchange couplings. Our results demonstrate that the TQC phase is highly sensitive to both the choice of coupling distribution and the values of the driving frequencies. These findings highlight promising experimental prospects for realizing TQCs in cold atomic systems and quantum simulators, providing valuable insights into their stability, dynamical properties, and potential for exploring novel non-equilibrium quantum phases.
@@ I Introduction
Time crystals (TCs) represent a fascinating phase of matter that emerges from the spontaneous breaking of time translational symmetry (TTS) in periodically driven quantum systems r1 ; r2 ; r3 ; r4 ; r5 ; r6 ; Yousefjani1 ; Yousefjani2 . These systems are typically subjected to external periodic perturbations, allowing them to maintain coherence and exhibit long-lived states characterized by discrete time translational symmetry breaking. In this context, preventing heating is crucial to avoid a transition to infinite temperature, which would lead to the loss of any temporal order r7 ; r8 .
@@
In disordered closed quantum systems, many-body localization (MBL) acts as a stabilizing mechanism against heating r99 ; r9 . MBL prevents the system from thermalizing by confining it to a localized region of the Hilbert space, thereby violating the eigenstate thermalization hypothesis (ETH). As a result, the quantum order is preserved over extended time scales, as the system retains its initial state within its eigenstates, shielded from thermalization processes r10 ; r11 ; r12 ; Pr12 ; rr12 ; rp12 .
@@
Periodically driven Floquet systems typically exhibit non-conservation of energy due to continuous energy absorption from external fields. However, the presence of Floquet MBL alters this dynamic. By breaking the ETH, Floquet MBL prevents the system from reaching thermal equilibrium, leading to the emergence of nonequilibrium phases, including discrete TC phases r13 ; eqr13 ; r14 ; r15 .
@@
In discrete TC phases, the dynamics of observables in the system are characterized by periodic responses that differ from the driving period. Specifically, for a Hamiltonian that oscillates with period $T$, the response functions of the system will exhibit oscillations with a period $nT$ (where $n\geq 2$) r8 .
Time quasicrystals (TQCs) represent a fascinating extension of the concept of time crystals (TCs). Unlike TCs, which break discrete time-translation symmetry by exhibiting a periodic response at a subharmonic frequency (a multiple of the driving frequency), TQCs display a more complex form of temporal order. They respond at multiple incommensurate frequencies, frequencies that are not integer multiples of the fundamental driving frequency, leading to quasiperiodic behavior in time.
@@
TQCs arise from quasiperiodic driving of a system where the time-dependent Hamiltonian $H(t)$ incorporates irrationally related frequencies. This leads to quasiperiodic dynamics, resulting in a non-repetitive evolution of the system’s state that densely explores a multi-dimensional toroidal phase space, without revisiting previous configurations. Such quasiperiodicity facilitates intricate dynamical behavior, including complex energy level spacings and dense trajectories in phase space rr17 ; dumitrescu2018 ; else2020 .
Recently, much attention has been devoted to the systems under quasiperiodic driving protocols verdeny2016 ; nandy2017 ; dumitrescu2018 ; giergiel2019 ; ray2019 ; zhao2021 ; pre13 ; else2020 ; mukherjee2020 ; FTQ ; FTS ; rr15 ; r17 ; r18 ; pre18 ; pre19 . In a recent proposal, it has been demonstrated that discrete TQCs can be realized in a Rydberg atomic chain by coupling two discrete TCs, where associated two external driving frequencies have the maximum incommensurability. In this system, the quasiperiodic response emerges due to interaction-induced synchronization between the subsystems pre24 .
It has been established that quasiperiodic driving ultimately leads to thermalization toward a featureless infinite-temperature state, but this process occurs after a long timescale known as the prethermal time dumitrescu2018 . This behavior is analogous to the prethermalization observed in high-frequency periodic driving, where the system first relaxes into a long-lived metastable state before eventual thermalization pre1 ; pre2 ; pre3 ; pre4 ; pre5 ; pre6 ; pre7 ; pre8 ; pre9 ; pre10 .
While rigorous results regarding heating rates are available for continuously varying quasiperiodic driving else2020 , there are also findings pertaining to discrete quasiperiodic driving protocols, such as those based on Fibonacci dumitrescu2018 and Thue-Morse sequences pre13 ; zhao2021 . Nonetheless, the general behavior under arbitrary quasiperiodic driving remains largely unexplored and requires further investigation.
The time evolution of entanglement entropy is a crucial parameter for distinguishing different dynamical regimes in driven many-body systems. In the MBL phase, the entanglement entropy grows logarithmically with time, $S(t)\sim\log t$, reflecting the slow dephasing dynamics and the absence of thermalization. In contrast, thermalizing systems exhibit a linear growth, $S(t)\sim t$, indicative of rapid entanglement spreading and eventual thermal equilibriumdumitrescu2018 ; pre20 . Under quasiperiodic driving, the entropy exhibits sublinear, power-law growth, $S(t)\sim t^{\alpha}$ with $\alpha<1$, characteristic of the prethermal regime and suggestive of a long-lived, non-thermal quasi-stationary state pre21 ; pre22 .
In this study, we investigate the disordered ITF model. Recent research has demonstrated that the clean ITF model can exhibit a discrete TC phase driven by global periodic kicks generated by a transverse magnetic field, without the need for Hamiltonian quenching or long-range interactions Yu2019 . Building on this, we analyze the dynamical properties of the disordered ITF chain under periodic magnetic driving. Our results reveal the presence of a robust discrete TC phase in the disordered system. We further assess the stability of this phase against external perturbations and imperfections in spin rotations, finding it to be remarkably resilient. Using exact diagonalization, we observe that increasing the chain length enhances the stability of the discrete TC.
Additionally, we explore a prethermal TQC phase within a quasiperiodically driven ITF model. By leveraging prethermalization mechanisms, we demonstrate the emergence of a long-lived phase characterized by magnetization oscillating at multiple incommensurate frequencies. This TQC phase exhibits strong coherence despite interactions and imperfections in the quasiperiodic driving, persisting over extended timescales before eventual thermalization. Furthermore, we examine the manifestation of TQCs in disordered ITF chains with exchange couplings randomly drawn from a symmetric distribution. In this context, the system can reside in a spin glass phase. By examining various distributions of the exchange couplings $J_{i}$ and different driving frequencies, we analyze the conditions under which the TQC appears in the spin glass phase.
@@
This paper is structured as follows: In Section II, we introduce the ITF spin chain, which is periodically driven by an external magnetic field, and analyze its magnetization in various directions, demonstrating the existence of a discrete TC phase. We further investigate the robustness of this discrete TC phase against perturbations in Section III. Section IV provides an overview of the definition and general properties of a TQC phase. In Section V, we incorporate a sinusoidal driving field into the disordered ITF chain, revealing the existence of a TQC phase and exploring its stability under mechanism of prethermalization. Section VI discusses the emergence of a TQC phase within an ITF chain exhibiting a spin glass phase. Finally, we summarize our findings in Section VII.
@@ II TC in disordered ITF model
$H(t)=\begin{cases}\hat{H}_{1}=g\sum\limits_{i}\hat{\sigma}_{i}^{x},\quad 0\leq t_{1}<t^{\prime}\\ \hat{H}_{2}=\sum\limits_{i}(-J_{i}\hat{\sigma}_{i}^{z}\hat{\sigma}_{i+1}^{z}+h_{i}\hat{\sigma}_{i}^{x}),\quad t^{\prime}<t_{2}\leq T\end{cases}$ (1)
Here, $t_{1}$ and $t_{2}$ are two dimensionless time scales satisfying $T=t_{1}+t_{2}$, where $T$ is a period indicating the time translational symmetry of the Hamiltonian, i.e. $\hat{H}(t)=\hat{H}(t+T)$. In Eq. (1), $g$ is a static transverse magnetic field (set to $g=\frac{\pi}{2}$), $J_{i}$ is site-dependent exchange coupling, randomly chosen from the interval of $\frac{1}{2}J,\frac{3}{2}J]$ with $J>0$, and $h_{i}$ is a site-dependent magnetic field, randomly chosen from the interval of [h,h][-h,h], where $h$ is an arbitrary positive value. The Hamiltonian $\hat{H}_{2}$ possesses the global $Z_{2}$ symmetry with the Ising parity operator $\hat{P}=\prod_{i=1}^{L}\hat{\sigma}_{i}^{x}$, which rotates all spins around the $x$ axisA berif .
In the clean system ($\delta J=J_{i}-J=0$ and $\delta h=h_{i}-h=0$), $\hat{H}_{2}$ reduces to the uniform ITF Hamiltonian. In this case, for $h\gg J$, the system is in the paramagnetic (PM) phase, where the ground state is gapped and nondegenerate, and all spins are aligned with the transverse field $h$, illustrated as $|\rightarrow\rightarrow\rightarrow\dots\rangle$. In this phase the ground state has no long-range order, the magnetization $m^{z}$ is vanishing, and spin-spin correlations decay exponentially by increasing spin-spin separation distance. As the exchange coupling strength $J$ is increased, it promotes the alignment of spins along the $z$ direction, leading to the breaking of Ising symmetry at the critical point $h=J$, where a phase transition occurs from the PM phase to a FM phase. In the FM phase, the ground state is characterized by either $|\uparrow\uparrow\dots\rangle$ or $|\downarrow\downarrow\dots\rangle$, resulting in a finite magnetization and a definitive breaking of the Ising symmetry.
@@
The situation is different for finite size chains. The system can tunnel between the two degenerate eigenstates at finite size, and consequently the two lowest lying eigenstates of the system are the following $Z_{2}$ symmetric states: $|\pm\rangle=\frac{1}{\sqrt{2}}[|\uparrow\uparrow\dots\uparrow\rangle\pm|\downarrow\downarrow\dots\downarrow\rangle]$. These states are the eigenstates of the Ising parity operator $\hat{P}$ with the eigenvalues $\pm 1$, they are long-range correlated with vanishing magnetization r7 .
@@
At finite temperatures, the Mermin-Wagner theorem asserts that one-dimensional systems with discrete symmetries cannot undergo spontaneous symmetry breaking, as thermal fluctuations in these systems are strong enough to disrupt long-range order, resulting in a finite density of excitationsmermin-wagner . In the FM phase, these excitations manifest as domain walls. The energy associated with these domain walls is typically independent of their position along the chain, allowing them to move freely. This mobility of domain walls contributes to the destabilization of any potential long-range order, thus reinforcing the conclusion that spontaneous symmetry breaking cannot be realized in one-dimensional systems at finite temperatures.
@@
The excited states can be described as superpositions of states featuring finite densities of domain walls distributed across various spatial locations. These delocalized and fluctuating domain walls introduce perturbations that disrupt long-range spin correlations at any finite temperature.
@@
The introduction of disorder, characterized by a non-clean regime with $h_{i}\in[-h,h]$ and $J_{i}\in[J/2,3J/2]$, leads to significant modifications in the system’s physical properties. In the absence of external magnetic fields, the system resides in a FM phase characterized by a non-zero magnetization $m^{z}$, as all exchange couplings are positive. Upon the application of a weak magnetic field, specifically when $h\ll J$, domain walls emerge along the chain. For a single domain wall positioned at $d$, the excited state of the system can be expressed as $|d\rangle=|\uparrow\uparrow\dots\uparrow\downarrow_{{}_{d}}\uparrow\uparrow\dots\rangle$. In contrast to the clean system, where a thermally activated domain wall can move freely, in the presence of disorder, the domain wall becomes immobilized due to the energy cost associated with its dispersion throughout the chain. As the magnetic field increases, the system can sustain a finite density of domain walls, all of which remain pinned by the disorder. It has been demonstrated that for $h<J$, the system exhibits characteristics of the many-body localized (MBL) phase within the FM phase, which remains stable in the presence of perturbations such as interactionsr23 .
@@
We now examine the potential for spontaneous breaking of time translational symmetry in the Floquet ITF model. The time evolution of the system, as described by equation (1), is governed by the Floquet operator given by:
@@
With $t_{1}=t_{2}=1$, the total driving period of the Hamiltonian is $T=2$, and accordingly, the drive frequency is $\omega=\frac{2\pi}{T}=\pi$. Also the Floquet operator is simplified to:
@@
Here, the first term captures the time evolution over the interval $1<t_{2}\leq T$ under the Hamiltonian $\hat{H}_{2}$, while the second term represents a spin rotation about the $x$-axis. Through this formalism, we aim to determine the conditions under which time translational symmetry may undergo spontaneous breaking in the context of the FITF model.
Consider the system initially in a fully polarized state along the $z$ axis, represented as $|\psi_{0}\rangle=\otimes_{i=1}^{L}|\uparrow_{i}\rangle$, where $L$ is the number of sites. The stroboscopic dynamics of the system are obtained by iteratively applying the Floquet time evolution operator, yielding the system’s state at discrete times $t=nT$, where $n$ denotes the number of driving periods, along with the time dependence of various order parameters. We analyze the temporal evolution of the order parameters, namely the magnetizations along the $x$, $y$, and $z$ axes, defined as
$m^{x,y,z}(t)=\left\langle\frac{1}{L}\sum_{i}^{L}\langle\psi(t)|\hat{\sigma}_{i}^{x,y,z}|\psi(t)\rangle\right\rangle,$ (4)
where $|\psi(t)\rangle$ represents the system’s state at time $t$, and $\langle\dots\rangle$ denotes the ensemble average over different random realizations.
In the presence of disorder, it is essential to compute ensemble averages of the magnetizations across various random realizations. By averaging over 150 random realizations, we derive the magnetizations for 160 Floquet cycles. Our findings indicate that when $\bar{J}\gg\bar{h}$ ($\bar{h}$ and $\bar{J}$ represent the mean values of the random fields $h_{i}$ and the random couplings $J_{i}$, respectively), the magnetization $m^{z}(t)$ exhibits oscillatory behavior around zero with a period of $2T$ (refer to Fig. 1). The periodic oscillations of $m^{z}(t)$, with a period that is double that of the Hamiltonian, suggest a breakdown of time translational symmetry. This phenomenon is indicative of discrete TC, provided that the observed oscillations remain stable in the thermodynamic limit ($L\rightarrow\infty$) and are resilient against perturbations such as interactions.
Figure 1: Time evolution of the magnetizations along $x,y$, and $z$ directions for a chain of 12 sites. For $J=5.5$ and $h=0.3$ (up), the magnetization along the $z$ axis exhibits periodic oscillations with a period of $2T$, while the $x$ and $y$ components remain zero. The inset plot illustrates the fast Fourier transform (FFT) of $m^{z}(t)$, highlighting a pronounced peak at $\omega=\pi/2$, which signifies the periodic nature of the magnetization. Conversely, for $J=0.3$ and $h=4.5$ (down), the magnetization oscillations are characterized by a lack of periodicity and stability. The peak value in the FFT is multiplied by 0.01, and this adjustment is applied to all other plots as well.
@@
III Stability of the discrete TC phase
In this section, we investigate the stability of the discrete TC in the presence of various external perturbations. We will demonstrate that the introduction of diverse perturbations, including interactions and imperfect rotations (deviations from perfect rotation), does not compromise the TC order.
III.1 Stability of the discrete TC phase under interactions
In this part, we examine the stability of the discrete TC phase under the influence of exchange interactions along the $x$-axis. In this context, the second term of the Hamiltonian in Eq. (1) is modified as follows: